Mitts Cosmetics Co.'s stock price is $55.94, and it recently paid a $1.00 dividend. This dividend is expected to grow by 29% for the next 3 years, then grow forever at a constant rate, g; and rs = 16%. At what constant rate is the stock expected to grow after Year 3? Round your answer to two decimal places.
Answer #1

Current Stock Price = P0 = $ 55.94, Current Dividend = $ 1, Initial Growth Rate = 29% for 3 Years, Let the perpetual growth rate be g

Required Return = 16 %

Initial Dividends: D1 = 1.29 x 1 = $ 1.29, D2 = 1.29 x 1.29 = $ 1.6641 and D3 = 1.66141 x 1.29 = $ 2,14669

Total Present Value of Initial Dividends = 1.29 / 1.16 + 1.6641 / (1.16)^(2) + 2.14669 / (1.16)^(3) = $ 3.724

Present Value of Perpetual Dividends = 55.94 - 3.724 = $52.216

Present Value of Perpetual Dividends = [2.14669 x (1+g) / (0.16-g)] x [1/(1.16)^(3)] = 52.216

2.14669 + 2.14669 g = 13.0406 - 81.5037 g

g = 0.13023 or 13.023 % ~ 13.02 %

> I got lost towards the end, where did the 13.0406 and the 81.5037 come from?

Meagan Thomas Sun, Nov 21, 2021 10:11 AM